Let S={2, 4, 7}

and T={1, 3, 5}

. Find f(S×T)

if  

  • f(x,y)=14x/3y
  •  


1
Expert's answer
2021-04-28T06:31:46-0400

Let "S=\{2, 4, 7\}" and "T=\{1, 3, 5\}", "f(x,y)=\frac{14x}{3y}."

Let us find "S\u00d7T:"

"S×T=\{(2,1),(2,3),(2,5),(4,1),(4,3),(4,5),(7,1),(7,3),(7,5)\}" .


Let us find "f(S\u00d7T):"


"f(2,1)=\frac{28}{3}, f(2,3)=\frac{28}{9}, f(2,5)=\frac{28}{15}, f(4,1)=\frac{56}{3}, f(4,3)=\frac{56}{9},"


"f(4,5)=\frac{56}{15}, f(7,1)=\frac{98}{3}, f(7,3)=\frac{98}{9}, f(7,5)=\frac{98}{15}."


It follows that "f(S×T)=\{\frac{28}{3}, \frac{28}{9}, \frac{28}{15}, \frac{56}{3}, \frac{56}{9},\frac{56}{15}, \frac{98}{3}, \frac{98}{9}, \frac{98}{15}\}."